Neutron stars, such as the one at the center of the Crab Nebula, have about the same mass as our sun but a much smaller diameter. Part A If you weigh 665 N on the earth, what would be your weight on the surface of a neutron star that has the same mass as our sun and a diameter of 16.0 km ? Take the mass of the sun to be ms = 1.99×1030 kg , the gravitational constant to be G = 6.67×10−11 N⋅m2/kg2 , and the acceleration due to gravity at the earth's surface to be g = 9.810 m/s2 . Express your weight wstar in newtons. View Available Hint(s) wstar w s t a r = nothing N Submit
Answer #1

Weight on the surface of neutron star will be given by:

W1 = m*g1

W1 = m*g*g1/g

W1 = W*g1/g

g1 = acceleration due to gravity at the neutron star's surface = ?

g = acceleration due to gravity at the earth's surface = 9.810 m/s^2

W = weight on earth = 665 N

Now

g1 = G*M1/R1^2

M1 = mass of neutron star = 1.99*10^30 kg

R1 = radius of neutron star = 16.0/2 = 8.0 km = 8000 m

So,

g1 = 6.67*10^-11*1.99*10^30/(8000^2)

g1 = 2073953125000 = 2.074*10^12 m/s^2

So,

Weight on neutron star will be

W1 = W*g1/g

W1 = 665*2.074*10^12/9.810

W1 = 1.406*10^14 N

Please Upvote.